A stock had returns of 17.48 percent, −4.92 percent, 20.15 percent, and 8.45 percent for the past four years. What is the variance of the returns?
Answer #1

Average return=Total return/Total time period

=(17.48-4.92+20.15+8.45)/4=10.29%

Return (Return-Average return)^2
17.48 (17.48-10.29)^2=51.6961
-4.92 (-4.92-10.29)^2=231.3441
20.15 (20.15-10.29)^2=97.2196
8.45 (8.45-10.29)^2=3.3856
Total=383.6454%

Standard deviation=[Total (Return-Average return)^2/(Time Period-1)]^(1/2)

=[383.6454/(4-1)]^(1/2)

=[383.6454/3]^(1/2)

=11.31%(Approx)

Variance=Standard deviation^2

=127.8818%